OpenStudy (anonymous):

please help i give medals! write a function rule for the table given below x | 0 | 1 | 2 | 3 | 4 | 5 | y |-3 | -2 | -1 | 0 | 1| 2 | a y=x+3 b y= -2x c y= x-3 d y= x+1

OpenStudy (anonymous):

select any two points as (0,-3) and (1,-2)

OpenStudy (anonymous):

then are you know what the method that used for form a straight line eqution bu using two points?

OpenStudy (anonymous):

no idea @esamalaa

OpenStudy (anonymous):

\[\frac{ y-y _{1} }{ x-x _{1} }=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

so would that be a, b, c, or d? D:

OpenStudy (anonymous):

@esamalaa ?? a b c or d?

OpenStudy (unklerhaukus):

x | 0 | 1 | 2 | 3 | 4 | 5 | y |-3 | -2 | -1 | 0 | 1 | 2 | can you see a common difference between x and y?

OpenStudy (anonymous):

@unklerhaukus i mean i guess top is positive bottom is negative im not sure to be honest

OpenStudy (unklerhaukus):

both x and y are increasing, right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

when x increases by one each time, how much does y increase?

OpenStudy (anonymous):

by -1

OpenStudy (unklerhaukus):

are you sure about that?

OpenStudy (anonymous):

OHH i think i get it now they both are increasing by 1, so the awnser would be d y=x+1 ?

OpenStudy (unklerhaukus):

they are both increasing by the same amount (one) each time, but they start at different values

OpenStudy (unklerhaukus):

what is the difference between the starting values?

OpenStudy (anonymous):

3 or -3

OpenStudy (unklerhaukus):

yeah -3

OpenStudy (unklerhaukus):

in fact the only difference between x and y is that y is three less than x , right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

Which of the options corresponds with the statement `y is three less than x` a) y=x+3 b) y= -2x c) y= x-3 d) y= x+1

OpenStudy (anonymous):

ah i think i see so it would be c x-3 for three less than?

OpenStudy (unklerhaukus):

you can try that solution out y= x-3 when x =0 y= 0-3 y= -3 which fits the first column on the table x | 0 | y |-3 |

OpenStudy (unklerhaukus):

what about the second column, for x=1 y= x-3 = 1-3 = ? x | 0 | 1 | y |-3 | ? | does this equation work?

OpenStudy (anonymous):

|1| |-2| i beleive so if you do -3-1=-2?

OpenStudy (unklerhaukus):

good, lets lets one more point , just to be sure what do you get if you sub x=5 into y=x-3

OpenStudy (unklerhaukus):

*lets test

OpenStudy (anonymous):

|5| |2| -3-5=2

OpenStudy (anonymous):

thanks so much man!

OpenStudy (unklerhaukus):

careful * x=5 into y = x -3 y = 5 -3 = 2

OpenStudy (anonymous):

so the answer would be C

OpenStudy (unklerhaukus):

yers